OpenStudy (anonymous):

solve. then round answer to the nearest hundredth. log_5x=3

OpenStudy (anonymous):

@dpflan ?

OpenStudy (anonymous):

\[log_5x=3\]Is that the equation?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

\[log_ab=x\] OK, this means, the number you raise a to in order to obtain x is b. so \[a^x = b\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

would it be x = 5^3 = 125?

OpenStudy (anonymous):

\[log_5x=3\] Is \[5^3 = x\]... Yeah you got it

OpenStudy (anonymous):

what is 125 to the nearest..?

OpenStudy (anonymous):

\[125 = 125.00000...\]

OpenStudy (anonymous):

just like the last one ;)

OpenStudy (anonymous):

what would 35 be rounded to the hundredth? 35.000?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Actually, no, you need one less 0. Using the decimal system, the values to the right of the decimal are fractional amounts with respect the base for the system, which is 10 here. So the first place would be \[10^{-1}\] which is 1/10, the second place is \[10^{-2}\] which is \[\frac{1}{10^2}=\frac{1}{100}\] , this that is the "hundredths" place

OpenStudy (anonymous):

so just two points to the right would be to the nearest hundrdeth

OpenStudy (anonymous):

cool..it didnt have anything to do with the previous question..

OpenStudy (anonymous):

i just wanted to know what 35 rounded to the nearest hundredth would be

OpenStudy (anonymous):

It's actually kind of cool, you use any number as the base. So if you have 123.456, then you have \[1*10^3 + 2*10^1 + 2*10^0 + 4*10^{-1} + 5*10^{-2} + 6*10^{-3}\]

OpenStudy (anonymous):

At least in the decimal system

OpenStudy (anonymous):

is that for 35?

OpenStudy (anonymous):

No, 35 is 35.00